OpenStudy (anonymous):

Which has the highest vapor pressure? A. 25 mL of water at 283 K B. 10 mL of water at 298 K C. 50 mL of water at 293 K D. 5 mL of water at 323 K

OpenStudy (anonymous):

Use formula PV/ T to find P for each one. U will know which one is higher.

OpenStudy (anonymous):

My highest number is my highest vapor pressure?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I'm sorry..just want to be sure so for example A) 25/283=0.883392226?

OpenStudy (anonymous):

b) 10/298=.033557047?

OpenStudy (anonymous):

am i doing it right?

OpenStudy (anonymous):

c) 50/293=.17064846 D) 5/323= .01547987 is my answer A?

OpenStudy (anonymous):

Wait min. I have another way we didnt need to calculate

OpenStudy (anonymous):

According Charles's law. Pressure is directly proportional to temperature. Now, u just look at which one have higher Temperature, u will have higher pressure

OpenStudy (anonymous):

Iam so sorry, the first one is wrong

OpenStudy (anonymous):

The second one, it will new right. Iam sure

OpenStudy (anonymous):

so D? has higher temperature..will D be the correct answer?

OpenStudy (anonymous):

So D is ur answers

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

for sure?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Remember. P, T will inversely proportional V. P is increasing T will be increase,and V desceae se